c) Determine the equilibrium price and quantity of the above market when j=100.

Qd = Qs given the following equations: 400 - 2P + 4j = - 100 + 3P and 400+400+100 = 3P+2P calculate price and quantity.


1
Expert's answer
2021-08-23T13:18:23-0400

"400-2P+4j=-100+3P"

"400-2P+4(100)=-100+3P"

"900=5P"

"P=180"

"Q= -100+3(180)= 440"

Price=180

Quantity= 440